OpenStudy (anonymous):

Find the Period of f(x) = 4sin^2 x + 7cos^2 x

OpenStudy (anonymous):

2(1-cos2x) + 7/2 + 7cos2x/2 = 11/2 + 3/2 (cos2x)

OpenStudy (anonymous):

Period of Cosx = 2pie so 3/2 * 2pie/2 = 3pie/2 .........Is it correct?

OpenStudy (anonymous):

@sauravshakya @satellite73

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

i think you have to take the LCM of the periods of these two

OpenStudy (anonymous):

But it is cos^2 x and sin^2 x

OpenStudy (ghazi):

oh i thought it's cos 2x and sin 2x

OpenStudy (ghazi):

i don't see any fault in your answer